# Scaling of Stochastic Normalizing Flows in $\mathrm{SU}(3)$ lattice gauge theory

### v1

@@ Appendix A Additional results and scaling in $\beta$
Appendix B Work distributions
@@
Figure 6: Distribution of the work $W$ (top panel) and of the exponential of the dissipated work $W-\Delta F$ (bottom panel) for $L/a=12$ for flows performed with NE-MCMC between $\beta=6.02$ and $\beta=6.178$ for 3 different values of $n_{\mathrm{step}}$. The black vertical line in the top panel represents the value of $\Delta F$ for this particular transformation in $\beta$.
Furthermore, the distribution of $\exp(-W)$ is also crucial, as it provides an immediate understanding of the quality of the sampling using eq. (11). In particular, in the bottom panel of fig. 6 we show how the values of $\exp(-W_{\rm d})$ from different evolutions behave for various values of $n_{\mathrm{step}}$ for a fixed transformation. While being further away for equilibrium (i.e. for smaller $n_{\mathrm{step}}$) the distribution is highly asymmetric, and the sampling with eq. (11) is reliable only if the tail on the right is populated enough, i.e. there are enough evolutions with $W<\Delta F$. On the other hand, when the transformation is closer to a reversible one (i.e. for the largest value of $n_{\mathrm{step}}$) the variance of the distribution is much smaller and less evolutions are needed for a correct estimation of observables and of the ratio of partition functions. This behavior is consistently described by the $\hat{\mathrm{ESS}}$, which precisely monitors the variance of this distribution, as understood in eq. (20).
@@ Acknowledgements.
We thank M. Albergo, S. Bacchio, C. Bonanno, M. Caselle, L. Giusti, G. Kanwar, P. Kessel, S. Nakajima, K. Nicoli, M. Panero, D. Panfalone, A. Patella, S. Schaefer, A. Tomiya, D. Vadacchino, L. Vaitl and L. Verzichelli for insightful comments and discussions. The numerical simulations were run on machines of the Consorzio Interuniversitario per il Calcolo Automatico dell’Italia Nord Orientale (CINECA). E. Cellini and A. Nada acknowledge support and A. Bulgarelli acknowledges partial support by the Simons Foundation grant 994300 (Simons Collaboration on Confinement and QCD Strings). A. Nada acknowledges support from the European Union - Next Generation EU, Mission 4 Component 1, CUP D53D23002970006, under the Italian PRIN “Progetti di Ricerca di Rilevante Interesse Nazionale – Bando 2022” prot. 2022TJFCYB. All the authors acknowledge support from the SFT Scientific Initiative of INFN.

### v2

@@ Appendix A Additional results and scaling in $\beta$
Appendix B Work distributions and the free energy
Understanding the behavior of the average over non-equilibrium evolutions that we perform, for example, when applying Jarzynski’s equality or the reweighting formula of eq. (11) is instrumental to obtain reliable results with this approach. The free-energy $\Delta F$ obtained directly from eq. (13) is the most natural candidate, as it has to be the same irrespective of the architecture being used. Results reported in fig. 6 have been obtained both with NE-MCMC and SNFs for a variable amount of $n_{\mathrm{step}}$ and show a general good agreement for both lattice sizes under study. Even if the computational effort is similar among the different architectures, flows with smaller $n_{\mathrm{step}}$ are less efficient and generally lead to larger errors. Furthermore, flows characterized by large $\tilde{D}_{\mathrm{KL}}$ or small $\mathrm{ESS}$ generally are at risk of overestimating $\Delta F$ if the number of evolutions is too low: as it will be clear in the following, this exponential average is susceptible to the presence of “rare events”.
Figure 6: Results for the dimensionless free energy $\Delta F$ for a flow between $\beta=6.02$ and $\beta=6.178$, for $L/a=12$ and $L/a=16$, for NE-MCMC (diamonds) and SNF (circles) architectures.
@@ Appendix B Work distributions and the free energy
Figure 7: Distribution of the work $W$ (top panel) and of the exponential of the dissipated work $W-\Delta F$ (bottom panel) for $L/a=12$ for flows performed with NE-MCMC between $\beta=6.02$ and $\beta=6.178$ for 3 different values of $n_{\mathrm{step}}$. The black vertical line in the top panel represents the value of $\Delta F$ for this particular transformation in $\beta$, chosen to be the most precise value for NE-MCMC from fig. 6. The sum of the histogram bars equals unity in all cases.
Furthermore, the distribution of $\exp(-W)$ is also crucial, as it provides an immediate understanding of the quality of the sampling using eq. (11). In particular, in the bottom panel of fig. 7 we show how the values of $\exp(-W_{\rm d})$ from different evolutions behave for various values of $n_{\mathrm{step}}$ for a fixed transformation. While being further away for equilibrium (i.e., for smaller $n_{\mathrm{step}}$) the distribution is highly asymmetric. This means that the estimation of $\Delta F$ with eq. (13) and the sampling with eq. (11) are reliable only if the tail on the right is populated enough, i.e., there are enough evolutions with $W<\Delta F$ (the “rare events” mentioned at the beginning of this section). On the other hand, when the transformation is closer to a reversible one (i.e., for the largest value of $n_{\mathrm{step}}$) the variance of the distribution is much smaller and less evolutions are needed for a correct estimation of observables and of the ratio of partition functions. This behavior is consistently described by the $\hat{\mathrm{ESS}}$, which precisely monitors the variance of this distribution, as understood in eq. (20).
@@ Acknowledgements.
We thank M. Albergo, S. Bacchio, C. Bonanno, M. Caselle, L. Giusti, G. Kanwar, P. Kessel, S. Nakajima, K. Nicoli, M. Panero, D. Panfalone, A. Patella, S. Schaefer, A. Tomiya, D. Vadacchino, L. Vaitl and L. Verzichelli for insightful comments and discussions. The numerical simulations were run on machines of the Consorzio Interuniversitario per il Calcolo Automatico dell’Italia Nord Orientale (CINECA). E. Cellini and A. Nada acknowledge support and A. Bulgarelli acknowledges partial support by the Simons Foundation grant 994300 (Simons Collaboration on Confinement and QCD Strings). A. Nada acknowledges support from the European Union - Next Generation EU, Mission 4 Component 1, CUP D53D23002970006, under the Italian PRIN “Progetti di Ricerca di Rilevante Interesse Nazionale – Bando 2022” prot. 2022ZTPK4E. All the authors acknowledge support from the SFT Scientific Initiative of INFN.